The sum of three numbers is 3. The first, minus the second, plus three times the third, is -7. The third plus 3 times the first, plus the second is 11. What are the numbers?
Answer by mananth(16946)

let the numbers be x , y , z.
x+y+z=3
z= 3-(x+y)
..
x+y-3z=-7
z+3x+y=11
..
plug the value of z in both the equations
x+y-3(3-(x+y)=-7
x+y-3(3-x-y)=-7
x+y-9+3x+3y=-7
4x+4y=2
2x+2y=1
..
z+3x+y=11
3-x-y+3x+y=11
2x=8
x=4
..
2x+2y=1
8+2y=1
2y=-7
y=-7/2
..
z=3-x-y
z=3-4+7/2
6-8+7 /2
=5/2
z=5/2